,f。he College outeur fur 3. True or False: In one-way ANOVA, all the samples must be the same size. 4. Consider the following sample values derived from monthly sales of three salespeople (in $000) АВС 15 15 19 10 10 12 9 5 11 16 16 12 17 12 16 The F-ratio is 3.44 and the null hypothesis is Ho: μ μ.. μ c a) Compute the degrees of freedom b) Find the critical value at á .05 3,g d) Do you reject Ho at á .05? Do Ne your conclusionsuisticli
Answer #1

Ans:

3)False

4)

a)N=15

k=3

df(numerator)=3-1=2

df(denominator)=N-k=15-3=12

b)

Critical F value=FINV(0.05,2,12)=3.885

c)

As,F=3.44 is less than 3.885,we fail to reject null hypothesis.

d)

There is not sufficient evidence to conclude that there is difference among three means.